OpenStudy (anonymous):

9) A rectangle has a length that is 5 feet more than its width. The area of this rectangle is 84 square feet. What is the width and length of this rectangle?

OpenStudy (anonymous):

84 = lw l = w + 5 solve.

OpenStudy (anonymous):

l=w+5 A=84 keep in mind: A=l*w, then substitute w+5 for l: 84=w+5*w. Can you solve this?

OpenStudy (anonymous):

7x12?

OpenStudy (anonymous):

yes, 7 is the width and 12 is the length. Good job :)